An electron orbits in a 3.8 mT field with angular momentum 8.6×10−26kgm2/s.

What is the diameter of the orbit? (Express your answer to two significant figures and include the appropriate units.-> meters)
Answer #1

the electron makes an orbit in a magnetic Field of radius r = mv/eB
The
angular momentum of a particle in an orbit is L = mvr. So, multiplying the expression for the radius of the orbit by r gives
r^2 = mvr/eB = L/eB

r = (L/eB)^1/2

r = (8.6*10^-26/(1.602 * 10^-19*3.8*10^-3))^1/2 = 0.011885 m
or
r = 1.2 cm

DIameter = 2.4 cm or 0.02377 m